1.     A hiker determines the length of a lake by listening for the echo of her shout reflected by a cliff at the far end of the lake. She hears the echo 2.5 s after shouting. Estimate the length of the lake.


1
Expert's answer
2021-03-21T11:36:27-0400

The distance traveled by the sound can be found as follows:


"d=vt."

From the other hand, the total distance from the hiker to the cliff and back equals:


"d=2l."

Then, we get:


"2l=vt,""l=\dfrac{vt}{2}=\dfrac{343\ \dfrac{m}{s}\cdot2.5\ s}{2}=430\ m."